# Generative Human Geometry Distribution

### v2

@@ 7 Conclusions
In conclusion, we introduce the generative human geometry distribution, the first method that learns the distribution of geometry distributions, enabling ‘infinite’ sampling to synthesize high-fidelity human geometries with realistic, pose-aware clothing details. Using flow matching, we learn a flow that transforms from the SMPL template to the target geometry, significantly improving training efficiency. Following this, our method encodes geometry into compact feature maps, facilitating various downstream generative tasks. Experimental results demonstrate that our approach achieves state-of-the-art quantitative performance in the pose-conditioned generation task compared to existing generative methods. Furthermore, in novel pose generation experiments, our method successfully synthesizes fine-grained clothing details that conform to human poses and exhibits strong robustness, generating plausible results even when provided with a feature map mismatched to the target pose. These results highlight the effectiveness of our representation and its potential for advancing 3D human modeling and synthesis.

### v3

@@ 7 Conclusions
In conclusion, we introduce the generative human geometry distribution, the first method that integrates geometry distributions into generative modeling. This distribution-over-distribution design is novel, unique, and particularly effective, making our approach the only one capable of producing fine-grained geometry in a generative framework. Specifically, using flow matching, we learn a flow that transforms from the SMPL template to the target geometry, significantly improving training efficiency. Following this, our method encodes geometry into compact feature maps, facilitating various downstream generative tasks. Experimental results demonstrate that our approach achieves state-of-the-art quantitative performance compared to existing generative methods. Furthermore, our method is able to synthesize fine-grained clothing details that conform to human poses and exhibits strong robustness, generating plausible results even when provided with a feature map mismatched to the target pose. These results highlight the effectiveness of our representation and its potential for advancing 3D human modeling and synthesis.
Appendix
Figure 10: Details of our auto-decoder. (a) We sample noise from a Gaussian distribution ($\mathbb{R}^{3}$) and generate the flow that transforms from the SMPL template distribution to the human geometry distribution. (b) A convolutional network is used to interpret and decompress the compact feature map with the SMPL vertex map. (c) The architecture details of our model.
Implementation details
Our training requires an infinite set of training pairs $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})$. In practice, we sample $2^{18}$ surface points $\mathbf{x}_{1}$ per human geometry. We observe that many points in loose clothing regions share the same nearest SMPL template location, requiring multiple samples from that location to reconstruct these regions. To alleviate this, we sample a slightly sparser set of $2^{17}$ points on the SMPL template and apply the k-nearest neighbors (KNN) algorithm to associate each $\mathbf{x}_{1}$ with the nearest template point. This distributes the mapping workload across different template locations, improving the sampling efficiency for loose clothing regions. During training, we randomly sample $30,000$ pairs per iteration.
As shown in Fig. 10, the compact feature map $z_{\mathcal{T}|\mathcal{S}}$ is set to $\mathbb{R}^{8\times 24\times 24}$. We adopt the linear and convolutional layers from EDM2 Karras et al. (2024; 2025), using $256$ channels in all layers. For optimization, we use AdamW Loshchilov (2017) with a learning rate of $8\times 10^{-3}$, following a cosine learning rate scheduler across all models. We apply L2 regularization with weight $1e^{-4}$ to the feature map during training.
Figure 11: The evolution of the Chamfer distance of training different distribution formulations.
We train both the auto-decoder and the generation model on the THuman2 dataset Yu et al. (2021b) for the pose-conditioned generation task, following previous methods Xu et al. (2022); Zhang et al. (2024d) to ensure fair comparisons. The auto-decoder model is trained for two days and the generation model for one day on a machine with two A100 GPUs. To enhance the model’s ability to generalize across diverse poses, we employ the classifier-guidance that randomly drops the pose condition with a probability of $0.5$. This encourages the network to learn the underlying data distribution rather than overfitting to specific poses. For the novel pose generation task, we utilize the 4DDress dataset Wang et al. (2024), which comprises 64 outfits across $520$ human motion sequences, totaling approximately $50,000$ human geometries with diverse poses and clothing styles. We train the auto-decoder on a machine with $4$ A100 GPUs for two days and train the novel pose generation model for one day. We don’t utilize the classifier-guidance strategy on this dataset as its larger scale naturally supports robust training.
The auto-decoder model has 56.0256.02M parameters. On an A100 GPU, the feature map decompression takes $81$ ms, which is executed only once for the entire diffusion process. Each denoising step requires $529$ ms for $1$ million points, with a peak memory usage of $9522$ MB. On memory-constrained devices, the model can sample multiple times with significantly fewer points per iteration. For instance, sampling $10,000$ points takes $71$ ms with a peak memory consumption of $1420$ MB.
We visualize our results by sampling $1$ million points for Gaussian Splatting (GS) Kerbl et al. (2023) rendering. Specifically, we render GS’s depth maps first, from which we then extract normal maps. All points share the same scaling value, determined by the average minimum distance between points.
Figure 12: Demonstrations of different methods with fixed geometry features across varying poses. All results show identical geometric details in different poses.
Figure 13: Novel pose generation results.
Geometry Distribution Formulation
For the experiment of different formulations in training on a dataset, the evolution of the Chamfer distance across $30,000$ iterations is illustrated in Fig. 11. Our proposed human geometry distribution achieves superior reconstruction quality, with a consistent decrease in the Chamfer distance.
Novel Pose Generation
Previous methods Zhang et al. (2023b); Xu et al. (2022); Zhang et al. (2024d) synthesize 3D humans in canonical space and obtain novel pose appearances via deformation. Their limitations are illustrated in Fig. 12, where we deform avatars to different poses, resulting in identical clothing details across all poses. In the main paper, we have showcased that our method can generate pose-aware deformations by synthesizing a pose-compatible feature map. Fig. 12 demonstrates that, although the wrinkle details may not be entirely natural, our method still produces reasonable results even when using an incompatible feature map.
Furthermore, we present additional novel pose generation results in Fig. 13. Specifically, we showcase challenging cases featuring intricate poses and complex clothing deformations, such as loose garments and skirts, which significantly deviate from the SMPL surface.
Use of Large Language Models
During the preparation of this manuscript, we employed a large language model (LLM) as an auxiliary tool to assist with grammar checking and minor language refinements. All scientific content, ideas, and claims were conceived and written by the authors. The final version of the manuscript was carefully reviewed, modified, and approved by the authors to ensure accuracy and clarity.